OpenStudy (anonymous):

Find the coordinates of the other endpoint when you are given the midpoint (point M) and one of the endpoints (point P). P = (3, 5) and M =(-2, 0)

OpenStudy (accessdenied):

Well, the midpoint formula gives you a relation between the endpoints and the midpoint. If you plug in the known information, you should be able to solve for the unknowns with that relation.

OpenStudy (accessdenied):

\[ M(x,y) = \left( \frac{x_1 + x_2}{2}, \frac{y_1 + y_2}{2} \right) \] If we plug in the known information: \[ M(-2, 0) = \left( \frac{3 + x_2}{2}, \frac{5 + y_2}{2} \right) \]

OpenStudy (anonymous):

thank you that helps

OpenStudy (accessdenied):

Then we can just relate the like-parts and solve: -2 = (3 + x)/2, 0 = (5 + y)/2

OpenStudy (accessdenied):

You're welcome!

OpenStudy (anonymous):

wait whats next i got -2=3x/2, 0=5y/2

OpenStudy (accessdenied):

You should have addition between 3x and 5y Other than that, we just solve for the x and y.

OpenStudy (accessdenied):

Like, for x: -2 = (3 + x)/2 Multiply everything by 2 -4 = 3 + x Subtract the 3 -7 = x We know x= -7 Same thing done for y.

OpenStudy (anonymous):

okay thank you